Solve the following problems. Show complete solution.


1. Find the volume of the parallel piped with the vertices in the given order: (0, 0, 0), (3, 0, 0), (0, 5, 1), (2, 0 , 5), (3, 5, 1), (5, 0, 5), (2, 5, 6), and (5, 5, 6).




1
Expert's answer
2020-03-19T16:12:37-0400

Consider a parallelepiped  the vertices of which have coordinates

"A(0,0,0)B(3,0,0), C(0,5,1), D(2,0,5), \\
A_1(3,5,1),B_1(5,0,5), C_1(2,5,6),D_1(5,5,6)" .

 Form vectors

"\overline{AB}=(3-0,0-0,0-0)=(3,0,0)\\
\overline{AC}=(0-0,5-0,1-0)=(0,5,1)\\
\overline{AD}=(2-0,0-0,5-0)=(2,0,5)"

"V=\overline{AB}\cdot (\overline{AC}\times \overline{AD})\\
\overline{AC}\times \overline{AD}=(\begin{vmatrix}
 5 & 1 \\
 0 & 5
\end{vmatrix},-\begin{vmatrix}
 0 & 1 \\
 2 & 5
\end{vmatrix},\begin{vmatrix}
0 & 5 \\
 2 &0
\end{vmatrix})=\\
=(25-0,-(0-2),0-10)=(25,2,-10)\\
V=3\cdot25+0\cdot2+0\cdot(-10)=75"

Answer: "V=75"